OpenStudy (anonymous):

OKAY...I realized my most recent question was randomly, easy sooo easy it was kinda stupid...soo get ready for a real brainbuster....... 2 x 2 = ?

OpenStudy (katrinakaif):

4

OpenStudy (kinggeorge):

I'm telling you, it's 42.

OpenStudy (bahrom7893):

if 2x2=? 2x=?/2 x=(?/2)/2 x=?/(2^2) x=?/4

OpenStudy (anonymous):

XD my questions excite people

OpenStudy (anonymous):

3.55555 rounded to 4 ;D

OpenStudy (anonymous):

King George has made a valid point...

OpenStudy (katrinakaif):

In his world, it seems quite valid.

OpenStudy (bahrom7893):

If 3 = 4, then 2*2 = 3

OpenStudy (anonymous):

yes young grasshopper...very valid in this lifetime hahaha